# Revisiting MAE pre-training for 3D medical image segmentation

### v1

@@ A.2 VolumeFusion
Volume Fusion (Wang et al., 2023) is a pseudo-segmentation task using two sub-volumes from different 3D scans, which are fused together based on random voxel-level fusion coefficients. The fused image is treated as input, and the model predicts the fusion category of each voxel, mimicking a segmentation task. Pretraining is optimized using a combination of Dice loss and cross-entropy loss.
@@
Volume Fusion has unique parameters defining the size ranges of the rectangles used for fusing together images. In our experiments we utilize a rectangle size range between [8, 100] sampled uniformly for each axis. This represent the 62.5% of our input patch size, and identical percentage as in the original paper. Moreover the amount of rectangles sampled is an important parameter. Like in the original paper we sample M Ũ(10, 40) different rectangles, iteratively. Lastly, the number of categories was chosen to be 5, as in the original paper (this represent K = 4).
@@ A.3 VoCo
The ’Volume Contrastive Learning Framework’ (VoCo) (Wu et al., 2024) is designed to enhance self-supervised learning for 3D medical image analysis by leveraging the consistent contextual positions of anatomical structures. The method involves generating base crops from different regions of 3D images and using these as class assignments. The framework then contrasts random sub-volume crops against these base crops, predicting their contextual positions using a contrastive learning approach. The authors utilize a Swin-UNETR model architecture, employing the AdamW optimizer with a cosine learning rate schedule for 100,000 pre-training steps. The specific hyperparameters include cropping non-overlapping volumes with a size of 64x64x64, and generating 4x4 base crops during the position prediction task. This represents an input patch size of 384×384×96 which is rescaled and resized to fit exactly 4x4 64x64x64 crops.
@@
Appendix B Additional results
Aside from the quantitative data on the development and test dataset, we provide the quantiative data of the ablation experiments here. The following additional results are provided: 1. Results when fine-tuning in a low-data regime are presented in Table 4. 2. Experiment on how to best transfer weights when transferring to a dataset with more than 1 input channel is provided in Table 6 3. Results on how the pre-training effects generalization is provided in Table 7. 4. Experiment results of investigating if one can reduce the fine-tuning steps are presented in Table 8
@@ Appendix B Additional results
Table 4: Forty images with SSL are almost as good as all data from-scratch! S3D-B model almost reaches the performance of the model trained from-scratch with only 40 training cases. * D1 has only 38 training cases for the train split.

### v2

@@ A.2 VolumeFusion
Volume Fusion [39] is a pseudo-segmentation task using two sub-volumes from different 3D scans, which are fused together based on random voxel-level fusion coefficients. The fused image is treated as input, and the model predicts the fusion category of each voxel, mimicking a segmentation task. Pretraining is optimized using a combination of Dice loss and cross-entropy loss.
@@
Volume Fusion has unique parameters defining the size ranges of the rectangles used for fusing together images. In our experiments we utilize a rectangle size range between [8, 100] sampled uniformly for each axis. This represent the 62.5% of our input patch size, and identical percentage as in the original paper. Moreover the amount of rectangles sampled is an important parameter. Like in the original paper we sample $M\sim\mathcal{U}(10,40)$ different rectangles, iteratively. Lastly, the number of categories was chosen to be 5, as in the original paper (this represent K = 4).
@@ A.3 VoCo
The ’Volume Contrastive Learning Framework’ (VoCo) [42] is designed to enhance self-supervised learning for 3D medical image analysis by leveraging the consistent contextual positions of anatomical structures. The method involves generating base crops from different regions of 3D images and using these as class assignments. The framework then contrasts random sub-volume crops against these base crops, predicting their contextual positions using a contrastive learning approach. The authors utilize a Swin-UNETR model architecture, employing the AdamW optimizer with a cosine learning rate schedule for 100,000 pre-training steps. The specific hyperparameters include cropping non-overlapping volumes with a size of 64x64x64, and generating 4x4 base crops during the position prediction task. This represents an input patch size of 384×384×96 which is rescaled and resized to fit exactly 4x4 64x64x64 crops.
@@
Appendix B Longer Training schedule
MAEs are known to benefit from increasing the length of the training schedule, as shown in He et al. [14]. We evaluate if this effect transfers to 3D medical pre-training by increasing the training batch size by x8 to $48$, the learning rate to $23e-$ 2, and the iteration steps by x5 to a total of $1.25M$ steps. We refer to this model as S3D-Long to denote the longer training schedule with more data seen. We denote that the architecture remains identical to the previous architecture, to isolate the effect of the length of the steps as well as the amount of samples seen. Results are presented in Tab. 12. It can be observed that this x32 actually leads to a decrease in overall model performance on our test datasets, showing a 0.6% lower average DSC as well as a 0.6% lower Average NSD.
The observed performance degradation of the S3D-Long model, despite the longer training schedule and increased data exposure, suggests several possible factors at play. While MAEs have shown benefits from extended training schedules in general computer vision tasks, the same assumptions may not directly transfer to 3D medical image pre-training due to the unique nature of this domain. The findings highlight the importance of tailoring training strategies to the domain.
Table 5: Public available checkpoint trained on the ABCD dataset: To provide a public available checkpoint, we retrained our proposed model on the ABCD dataset, indicated by a *. It performs slightly worse than the network pre-trained on the private dataset.
SSL Method No (Dyn.) No (Fix.) S3D* S3D
Dataset Dice Similarity Coefficient (DSC)
MS FLAIR (D1) 57.81 59.82 59.75 60.35
Brain Mets (D2) 63.66 56.53 64.20 65.24
Hippocampus (D3) 89.18 89.24 89.45 89.60
Atlas22 (D4) 63.28 65.52 66.61 66.95
CrossModa (D5) 85.64 83.44 83.61 84.08
Cosmos22 (D6) 60.28 78.17 80.01 80.00
ISLES22 (D7) 77.94 79.44 78.94 79.70
Hanseg (D8) 59.00 61.85 61.27 62.11
HNTS-MRG24 (D9) 66.73 65.90 67.03 68.62
BRATS24 Africa (D10) 93.07 92.51 92.49 92.19
Avg. DSC 71.66 73.24 74.34 74.88
Avg. Rank 3.2 2.9 2.4 1.5
Appendix C Additional results
Aside from the quantitative data on the development and test dataset, we provide the quantiative data of the ablation experiments here. The following additional results are provided: 1. Results when fine-tuning in a low-data regime are presented in Tab. 6. 2. Experiment on how to best transfer weights when transferring to a dataset with more than 1 input channel is provided in Tab. 8 3. Results on how the pre-training effects generalization is provided in Tab. 9. 4. Experiment results of investigating if one can reduce the fine-tuning steps are presented in Tab. 10
C.1 Public weights trained on the ABCD dataset
Due to patient privacy concerns and data ownership regulations, we are unable to share the original pre-trained weights. As an alternative, we retrained our best-performing model on the National Institute of Health’s Adolescent Brain Cognitive Development (ABCD) dataset. This dataset comprises about 41k MRI scans with a 50-to-50 ratio of T1-weighted to T2-weighted scans. The results, shown in Tab. 5, indicate that the original model slightly outperforms the version pre-trained on the ABCD dataset. This discrepancy is likely attributable to the greater diversity and variation in the images within our private dataset, which enables a more robust feature representation.
C.2 Comparison to previous work on CT data
Although this study focuses on brain MRI images, we also evaluate our method on a CT downstream task. Table 11 presents the results of our approach on the BTCV multi-organ segmentation task [20]. For comparison, we incorporate a diverse set of results reported in [30], which were trained using an unspecified 80/20 data split. Additionally, we fine-tune the publicly available HySpark checkpoint [30] using the same five-fold cross-validation split as for our method. Remarkably, M3D outperforms all other methods, despite being pretrained exclusively on brain MRI data. Notably, none of the related approaches surpass our backbone trained on scratch. This highlights the critical role of leveraging state-of-the-art networks and advanced training frameworks, such as nnU-Net. [18, 19].
Table 6: Forty images with SSL are almost as good as all data from-scratch! S3D model almost reaches the performance of the model trained from-scratch with only 40 training cases. * D1 has only 38 training cases for the train split.
@@ C.2 Comparison to previous work on CT data
Table 11: S3D outperforms all related work on the BTCV dataset [20]. Despite being pretrained exclusively on brain MRI data, our network outperforms all related methods. Values above the line are sourced from Tang et al. [30], which used an unknown 80/20 split. To fine-tuned their published pre-trained weights using a 5-fold cross-validation. Below the line, all models, including ours, were trained on the same 5-fold cross-validation. Notably, even though all related work leveraged CT data for pretraining, none surpassed the performance of our backbone model trained from scratch, emphasizing the importance of Pitfall 2.
Pre-training Method Network Spl Kid Gall Eso Liv Sto Aor IVC Veins Pan AG Avg
vox2vec [10] 3D UNet(FPN) [27] 91.40 90.70 59.50 72.70 96.30 83.20 91.30 83.90 69.20 73.90 65.20 79.50
SUP [12] Swin UNETR [12] 84.20 86.70 58.40 70.40 94.40 76.00 87.70 82.10 67.00 69.80 61.00 75.80
MAE [14] UNETR [13] 90.71 87.63 62.50 72.60 96.09 94.73 86.11 90.36 71.00 75.47 63.77 79.07
SimMIM [33] Swin UNETR [12] 88.33 86.82 62.43 74.36 92.35 90.70 83.03 87.43 68.04 68.43 58.65 76.44
SparK [33] MedNeXt [28] 90.92 87.66 62.43 74.36 95.03 84.85 86.04 80.63 68.83 76.57 61.43 79.21
HySparK [30] MedNeXt+ViT  [30] 90.67 88.32 68.18 74.20 95.03 87.46 90.17 84.50 70.04 78.36 66.75 80.67
No MedNeXt+ViT  [30] 90.35 87.46 63.18 74.49 95.09 86.00 89.29 83.22 71.85 79.48 62.59 80.27
HySparK [30] MedNeXt+ViT [30] 90.94 86.99 63.43 74.39 95.12 87.15 88.92 83.48 72.77 79.66 64.84 80.67
No (Dyn.) nnU-Net [18] 90.44 88.52 68.86 78.14 95.53 88.06 91.59 86.47 76.27 81.78 71.06 83.34
No (Fix.) ResEncL (fixed) [19] 91.97 89.58 68.76 79.18 95.96 91.97 92.80 87.16 77.29 84.01 72.21 84.63
S3D ResEncL (fixed) [19] 92.00 90.40 70.77 78.71 96.01 92.51 92.83 87.04 77.28 84.79 72.58 84.99
Table 12: Longer training schedule degrades performance. When training with a larger batch size, higher learning rate, and more train steps we observe a degradation in performance for DSC and NSD. Ranks are calculated only between the four methods presented in the table.
Dice Similarity Coefficient
Dataset No Dyn. No Fixed S3D S3D-Long
MS FLAIR (D1) 57.81 59.82 60.35 59.85
Brain Mets (D2) 63.66 56.53 65.24 64.81
Hippocampus (D3) 89.18 89.24 89.60 89.34
Atlas22 (D4) 63.28 65.52 66.95 64.58
CrossModa (D5) 85.64 83.44 84.08 84.02
Cosmos22 (D6) 60.28 78.17 80.00 80.01
ISLES22 (D7) 77.94 79.44 79.70 79.89
Hanseg (D8) 59.00 61.85 62.11 61.93
HNTS-MRG24 (D9) 66.73 65.90 68.62 67.94
BRATS24 Africa (D10) 93.07 92.51 92.19 92.90
T2 Aneurysms (D11) 46.76 41.97 47.26 44.15
Avg. DSC 69.40 70.40 72.37 71.77
Avg. Rank 3.09 3.27 1.55 2.09
Normalized Surface Distance
MS FLAIR (D1) 78.77 80.16 80.03 80.40
Brain Mets (D2) 80.72 76.72 82.53 82.32
Hippocampus (D3) 99.46 99.42 99.46 99.44
Atlas22 (D4) 70.52 73.77 75.35 73.45
CrossModa (D5) 99.85 99.76 99.81 99.80
Cosmos22 (D6) 72.60 96.47 97.45 96.75
ISLES22 (D7) 88.55 90.45 90.59 90.72
Hanseg (D8) 82.20 85.94 85.80 86.20
HNTS-MRG24 (D9) 71.83 71.26 74.07 73.17
BRATS24 Africa (D10) 95.66 95.36 95.06 95.72
T2 Aneurysms (D11) 62.24 55.56 61.18 57.07
Avg. NSD 82.04 84.08 85.58 85.00
Avg. NSD Rank 2.82 3.18 2.00 2.00
Figure 3: Distribution of our pre-training dataset. The dataset stems from 44 centers and includes 8400 Patients with a 60 to 40 female-to-male ratio. Most patients were imaged with a 1.5 Tesla Philips Achieva or Ingenia scanner. The most prevalent modalities are T1 and T2-weighted images with some additional FLAIR images present. While other modalities were in the dataset, these were not used as prevalence was deemed too low.